Air flows steadily through an engine at a constant temperature of 400k. Find the work per kilogram if the exit pressure is one-third the inlet pressure and the inlet pressure is 207 KPa. Assume that the kinetic energy and potential energy variation is negligible.


1
Expert's answer
2021-11-18T05:32:02-0500



Velocity of the ball at maximum height is zero, v=0

"v \n^2\n \u2212u \n^2\n =2aH"

"0\u2212(49) \n^2\n =2\u00d7(\u22129.8)\u00d7H"

H=122.5 m   


v=u+at

0=49−9.8t

t=5s

Total time taken by ball to return to the surface, T=2t=10 s